wiki:UC20170911

Version 1 (modified by trac, 16 months ago) ( diff )

--

APO 3.5-m Users Committee Phonecon, 9/11/17

Attending: Nancy Chanover, Mark Klaene, Rachel Kuzio de Naray, Mary Beth Kaiser, John Tobin, John Wisniewski, Jon Holtzman, Anne Verbiscer, Chip Kobulnicky, Bill Ketzeback, Sarah Tuttle, Ben Williams, Russet McMillan, Scott Anderson, Joanne Hughes Clark, Leslie Hebb, Kal Kadlec


These minutes were graciously provided by Anne Verbiscer. Thank you Anne!!


Kal Kadlec, a recently hired optical engineer in the University of Washington Telescope Engineering Group, joined us for his first Users Committee telecon. Welcome Kal!

User feedback and comments from institutional representatives:

  • Oklahoma - John T. had nothing new to report, had a good night of observing last night
  • Colorado – absent; no report
  • Johns Hopkins - Mary Beth had nothing new to report
  • Wyoming - Chip had nothing new to report
  • Washington - Scott had nothing new to report
  • NAPG - Leslie had nothing new to report
  • Seattle U - Joanne had nothing new to report
  • NMSU - Jon relayed that his users are glad to see the echelle back
  • Virginia – Anne asks that TripeSpec users refer to online users manual. Russet says it doesn’t answer the question of being able to tell if something is saturated. Nancy asks if level of detail is insufficient or whether it’s something else. Persistence shows up when you’re pretty sure that the previous images were not saturated. Nancy says they should iterate with UVA team so the users guide is as clear as possible.
  • Georgia State – Rachel had nothing new to report (winds from Irma just beginning to pick up)

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 8/13/17 through 9/06/17

0) Overview

Generally recovered from shutdown. Smoke from Montana fires has been with us for the last several days, more heavily in the valley below us. Rain has dried up to some extent and Hurricane Harvey failed to produce moisture in our area. We organized a partial eclipse event for the public at the Sunspot Astronomy Visitor’s Center that was attended by about 200 people.

1) Telescope

Telescope was operational with minimal issues. We had several issues linger as result of new terminal server communication protocol. We believe we have them all under control.

2) Instruments

DIS: Sarah and Bill warmed DIS cameras and serviced dewar. cleaned inside black contamination and adjusted focus shim in red camera. Installed CCD protected field lens holder in Red camera and new ion pump in blue. Field flattener lens was clear of contamination when opened. Note: we vacuum pumped cameras during warmup this time.

TripleSpec is operational.

Agile has been operational this period.

Echelle is operational. With significant support from at first Dale Sandford then Connie Rockosi, Ed Leon was able to track down a capacitor in the initialization circuit. In tracking this down Ed was able to get more complete documentation. The failure was only seen when the electronics were closed up, on the bench with the saddle bag open and using extender cards to trace signals it did not fail. This added significantly to the time required to track down the failed component. We are still chasing the apparent contamination on the detector despite thorough vacuum pumping this summer.

NIC-FPS is operational.

ARCTIC is operational with no change.

3) ARCSAT - reassembled, but MaximDL software is causing issues. While it is a known issue it currently is much worse making the telescope unusable. Due to recent echelle and DIS work we have not been able to work on this.


Additional telescope and instrument discussion:

The telescope is back to operation in its normal configuration coming out of shutdown. There is a new generation terminal server. The echelle is back and operational. Ed Leon found a failed capacitor in the initialization circuitry in the saddlebag electronics. This allowed us to do the troubleshooting downstairs in the instrument lab and get it back together. We do see an increase in inter-order light but we're not yet ready to say anything concrete. Other instruments are operating well. DIS went through a shutdown last week. Sarah serviced the dewar inside. Contamination appears to be coming from the ion pump. Field flattener optics did not show signs of water spots this time. Put in new flat field lens holder. Blue camera showed signs of problem with temperature sensor yesterday. Blue camera is back on the bench and pumping down. Red camera is cold. Bill had nothing to add. Optimistically DIS Blue should be cooled by tomorrow morning. DIS is scheduled for a half-night of engineering tomorrow night. To follow up on the ARCSAT summary given in Mark's report above, the primary mirror was recoated, but the coating is still not as good as we want. There appears to be some residue on the glass that cannot be removed. The decision was made to put it back in the telescope with the current coating. Still not comfortable with turning the telescope over for operations due to ongoing problems with the control software (MaximDL), which keeps crashing. We hope to make progress this coming week on ARCSAT.

Bill recapped the summer shutdown, which was relatively short this year (roughly 2 weeks). They realuminized the tertiary mirror. Due to the issues at the illuminizing chamber we put off doing the secondary for another year until we are confident that things are settled over there. Did some minor things to the telescope: cleaning, servicing of components that typically get dirty. Overall a pretty unremarkable shutdown. Servicing UPS batteries was major task for this summer.


Q3 3.5-m scheduling:

Everything is fine.


Q4 3.5-m scheduling:

Russet just published the Q4 schedule; it is linked through the web page and there will be an email announcement later today. October A-halves were oversubscribed, which is common. Some people are chasing setting targets due to bad weather earlier in the summer. The oversubscription is worse this year because programs that requested a lot of October A halves were the ones who lost echelle time in July, therefore not everyone's requests may have been optimally scheduled. We have a returning program from Georgia State, in collaboration with NAPG members, which is using the first hour of almost every night in November and December. Observers had a lot of problems with sacrificing the first hour when this was done previously (a couple of years ago) in the month of May, which had shorter nights. This should be better with longer nights this time around but there are some programs that will need twilight flats, etc. This program uses DIS. There will be some difficulty getting calibration for everyone. If you have an A half program scheduled in November or December, talk with the observing specialists about this. Russet won’t ask ob-specs to come in early for cals when shifts are 16 hours long. This could be a big problem with classes coming in. Question about how this scheduling decision was made, and how the observatory set its priorities? Russet says that the program is not unprecedented and this kind of scheduling flexibility was one of the reasons that Georgia State became part of ARC. Suzanne supported the previous observing campaign; we have new directors now who also support it. Rachel reiterated that this is the only observatory where a program like this works. Sarah requested more information regarding how this adversely affects a class trip. Answer is about getting twilight flats, etc. Question about doing this program during B half twilight, but there is interest in that time too. If this is a good science program then it may be a good use of the time. Nancy says it should be tried to see if it is a good use of the time. Need some information about it to see if it should be continued or redone in the future. If institutions have issues, they should go through local scheduler. There is also a lot of OPEN time available and folks should jump on it.


NA2 Baffle Project

Discussion within the engineering group about thoughts on redesigning the baffle for the secondary mirror so the full ARCTIC FOV can be used without vignetting. Nancy will circulate a document that describes our thoughts on how to proceed with study to analyze scattered light in the system. User reps should share with their institutional members.


APOGEE Fiber Feed Project

Trench was dug between 3.5m and the SDSS telescope. Nancy suggests tabling this for the time being. Comes down to manpower at this point because we’ve lost a couple of engineers, including the one who was the primary on this. Unless someone steps forward to take this on, it should be tabled for now. If we were to put it on NA1, the costs could be reasonable, but still very expensive.


TDAWG Update

The first meeting of Time Domain Astronomy Working Group was held a couple of weeks ago. The path is set out and website is built to take in ideas. Want to focus on what ARC institutions want to do rather than what NSF wants to do to follow up on LSST discoveries. Will be forming ideas and putting together a document. If people are interested in joining they still can do so. Will have telecons every two weeks. Part of this is learning how to use alert streams. Stuff is happening right now that will affect the future success of APO.


ARCSAT Usage Survey Results

Nancy sent the survey results out before the call. Got 16 responses, not an overwhelming number but OK. Pretty interesting read. Some isn’t so surprising. Some are using it for students for training. Pleasantly surprised to see that data have been reduced, analyzed and published. Range in comments. Some say it’s useless, others say it’s very useful. Questions were not designed to make linkages between those who want to use it for their own science and those who left the positive feedback. People who like it, use it, people who don’t like it stopped using it. A TUI-like interface would be useful. The current interface is a little funny although not too difficult to use. The primary problems with it are that it is unstable (keeps crashing!) and that it makes too many decisions for the observer. Nancy wanted to make sure all Committee members have a copy of the survey responses to keep for future reference.


Request for Science Highlights for ARC Board of Governors Meeting

This topic was discussed in June, when we asked is it useful to present these highlights at this meeting? It is potentially of use for the administrative board members to see the return on investment. Nancy would like to collect science highlights. Previously Suzanne had asked committee members to send them to her. Deadline is Friday, September 29. One Powerpoint slide per program on results from the past year. Historically we had given users until the second week in October; Mark worries that Sept. 29 is going to be rather quick for some people. We will see how many we get by 9/29 and then may need to issue a second deadline. Nancy would like extra time to digest them and synthesize them into a coherent story for the Board presentation. She doesn’t mind if users send them directly to her.


ACTION ITEMS:

open from previous meetings:

  • Nancy will work with Ben and Russet on a web page that lists the telescope scheduling policies to avoid any future misunderstandings. Status: OPEN. Nancy sent out an email to the institutional schedulers, which then generated some questions that need to be cleared up. A new email and a web page outlining these policies still need to be developed. No change from last month; the scheduling team has been overtaken by the echelle events and hasn't had time to revisit this. Nancy promised to get this completed over the next month!
  • Users Committee members should distribute the AS4 overview and joining document and poll their users about interests in AS4. Status: CLOSED. Nancy sent out the document last night; UC reps should distribute it to their departments and discuss whether there scientific interest in getting involved in AS4, either at the individual department level or perhaps at the ARC level.

new from this meeting:

  • Russet will iterate with the UVA team to clarify some of the language in the TripleSpec User's Guide.
  • UC reps should request science highlights from their department members - one Powerpoint slide per program - for sharing at the ARC Board of Governors meeting in early November. The slides can be sent directly to Nancy or collected by the UC reps and sent to Nancy as one packet.
  • Nancy will distribute a document to the UC reps concerning thoughts on the NA2 baffle project.

Next meeting:

The next Users Committee will be on Monday 09 October at 11:00 am MDT.

Note: See TracWiki for help on using the wiki.